OpenStudy (anonymous):

Find g(f(x)) given: f(x) = 2x – 8 and g(x) = -5x – 10

OpenStudy (anonymous):

g(f(x)=-5(2x-8)-10 =-10x+40-10 =-10x+30 -30=-10x x=3

OpenStudy (anonymous):

g(f(x)) = -5(2x-8)-10 =-10x+40-10 =-10x+30